# EasyRec: Simple yet Effective Language Models for Recommendation

### v3

@@ 4.2. Performance Comparision for Text-based Recommendation (RQ1)
4.2.1. Baseline Methods and Settings
For our comparative evaluation, we included a diverse set of language models as text embedders: (i) General Language Models: BERT (Devlin et al., 2018), RoBERTa (Liu et al., 2019), and BART (Lewis et al., 2019); (ii) Language Models for Dense Retrieval: SimCSE (Gao et al., 2021), GTR (Ni et al., 2021), and BGE (Xiao et al., 2023); (iii) Pre-trained Language Models for Recommendation: BLaIR (Hou et al., 2024). Additionally, we also compared against the state-of-the-art text embedding models provided by OpenAI. The detailed descriptions of these baseline models can be found in Appendix A.3. This selection of language models covers a broad spectrum, from general-purpose contextual encoders to specialized models tailored for tasks like dense retrieval and recommendation.
4.2.2. Result Analysis
The overall comparison of different models is presented in Table 3. This evaluation reveals several noteworthy observations, which are outlined below:
  • Superiority across Diverse Datasets. Our evaluation consistently shows that the EasyRec outperforms all other models across the three datasets spanning different platforms. This provides strong evidence for the effectiveness of the EasyRec. We attribute these improvements to two key factors: i) By injecting collaborative signals into the language models, we effectively optimized our EasyRec using supervised contrastive learning within the recommendation context. This approach allows the model to inherently encode user and item text embeddings that are well-suited for recommendation tasks. ii) By integrating a diverse array of datasets across multiple categories and utilizing data augmentation techniques to enrich the text descriptions for training, our EasyRec exhibits impressive generalization capabilities, enabling it to effectively handle unseen data.
  • Scaling Law Investigation of EasyRec Model. Our experiments revealed that as the size of the EasyRec model increases (from small to large), its performance consistently improves across all three datasets. This observation reflects a scaling law, where the model’s performance growth is directly correlated with its size. Furthermore, this finding effectively reinforces the validity of text-based recommendation systems. It also validates our approach to training the language model, which enables it to learn collaborative signals from a new perspective.
4.2.3. Impact of Training Objectives
Furthermore, we evaluate the impact of different training objectives on the language model’s learning process. To this end, we implemented EasyRec-Large training using BPR loss (i.e., one negative item per training sample) for comparison with the contrastive learning results. This approach allows us to directly assess how the choice of training objective influences model performance. As shown in Table 4, the performance of the model trained with contrastive learning generally outperforms that of the model trained with BPR loss. This outcome highlights the effectiveness of employing contrastive learning to better incorporate collaborative information into the language models, thereby enhancing their overall performance in recommendation tasks.
4.3. Performance of Text-enhanced CF (RQ2)
Table 5. Recommendation performance in text-enhanced collaborative filtering. The experiment was conducted on the Steam dataset with 5-runs to obtain the mean results.
@@ 4.3. Performance of Text-enhanced CF (RQ2)
BERT 0.0500 0.0822 0.1313 0.0556 0.0663 0.0829
RoBERTa 0.0517 0.0848 0.1351 0.0573 0.0684 0.0854
BART 0.0529 0.0874 0.1383 0.0585 0.0701 0.0874
SimCSE 0.0529 0.0877 0.1395 0.0588 0.0706 0.0881
BLaIR 0.0535 0.0880 0.1392 0.0593 0.0708 0.0882
GTR 0.0532 0.0873 0.1387 0.0592 0.0706 0.0880
BGE 0.0535 0.0875 0.1393 0.0591 0.0705 0.0881
EasyRec 0.0540 0.0881 0.1402 0.0597 0.0712 0.0888
@@
BERT 0.0518 0.0849 0.1347 0.0575 0.0684 0.0852
RoBERTa 0.0531 0.0867 0.1374 0.0587 0.0699 0.0870
BART 0.0540 0.0887 0.1407 0.0599 0.0715 0.0891
SimCSE 0.0541 0.0891 0.1417 0.0602 0.0719 0.0898
BLaIR 0.0551 0.0897 0.1418 0.0609 0.0724 0.0901
GTR 0.0542 0.0894 0.1417 0.0601 0.0719 0.0896
BGE 0.0547 0.0891 0.1407 0.0603 0.0718 0.0893
EasyRec 0.0554 0.0908 0.1430 0.0614 0.0732 0.0908
In addition to our investigation of zero-shot recommendation scenarios, we explore the potential of EasyRec as an enhancement when integrated with CF models. To assess the effectiveness of various LMs in CF, we employ two widely used ID-based methods as backbone models: GCCF (Chen et al., 2020) and LightGCN (He et al., 2020), which were chosen for their proven effectiveness and efficiency. Furthermore, we utilize the advanced model-agnostic text-enhanced framework RLMRec (Ren et al., 2024a) with contrastive alignment to conduct our investigation. We compare the large versions of both EasyRec and other open-source LMs. The key findings from the results in Table 5 are:
  • Compared to the backbone methods alone, the integration of the text-enhanced framework generally improves the performance for both GCCF and LightGCN. This observation highlights the significance of incorporating text modality into the recommendation paradigm, as it can enhance user and item profiles.
  • Among the various LMs tested, EasyRec consistently achieves the highest performance in the text-enhanced recommenders. This outcome not only illustrates the efficacy of EasyRec for recommendation tasks, but also emphasizes the advantages of incorporating collaborative information into language models.
4.4. Effectiveness of Profile Diversification (RQ3)
@@
Figure 4. Performance with respect to data size. ”Augmentation Count” indicates the number $t$ of diversified profiles.
@@
  • Effectiveness of Profile Diversification. The increase in the number of diversified profiles (from 0 to 3) enhances model performance, particularly for larger models. This finding underscores the effectiveness of our augmentation approach using LLMs for profile diversification, and emphasizes the significance of increasing training data for improved outcomes.
  • Scaling Relationship: The scaling experiments on both model size and data size reveal a crucial relationship that influences model performance. This demonstrates that our approach of training the language model with collaborative signals follows a scaling law, indicating that model performance benefits from both increased capacity and data volume. Such scaling laws are vital as they provide insights into how model capacity and data availability interact, guiding future research and development.
4.5. Model Fast Adaptation Case Study (RQ4)
Figure 5. Case study on handling user preference shift.
@@ 4.5. Model Fast Adaptation Case Study (RQ4)
We visualize all the encoded embeddings using t-SNE (Van der Maaten and Hinton, 2008), which reveals a significant shift in the user embeddings within the feature space. Correspondingly, recommended items transition from basketball-related products to swimming gear, reflecting the user’s changing preferences. Notably, this adjustment is accomplished solely by modifying the user’s profile, without the need for further training of the model. This underscores the efficiency and flexibility of our approach in adapting to evolving user preferences.
5. Related Work
LMs-Powered Recommender Systems. Recent advancements in recommender systems increasingly incorporate textual modalities (Yuan et al., 2023), thus enhancing traditional approaches. The semantic representations encoded by pre-trained language models are essential features for improving recommender models, particularly in click-through rate prediction (Xi et al., 2023; Geng et al., 2024) and transferable sequence recommendations (Hou et al., 2022). These embeddings capture informative content relevant to recommendation tasks (Sheng et al., 2024). Some works also leverage text-based agents to enhance performance (Zhang et al., 2024a, b). A notable recent contribution is RLMRec (Ren et al., 2024a), a text-enhanced framework that improves ID-based recommenders using principles from information theory. However, many prior studies have relied on general text embeddings, such as BERT-based models (Ding et al., 2021; Yuan et al., 2023) or proprietary OpenAI embeddings (Ren et al., 2024a), rather than those specifically tailored for recommendation purposes. A recent work BLaIR (Hou et al., 2024) leverages the item metadata and interaction-level user feedback on this item for LMs training, yielding promising results in query-based item retrieval. In contrast, our approach assigns each user and item a collaboratively generated profile that reflects their preferences. EasyRec optimizes the learned correlations between user and item entities using CF signals, which not only demonstrates impressive zero-shot performance but also enhances text-augmented results.
@@ 5. Related Work
Cross-Domain Recommendation. The fundamental concept behind cross-domain recommendation is to enhance recommendations in one domain by leveraging data from another domain, which is typically more abundant, to address data sparsity and improve personalization (Zang et al., 2022). For instance, the graph collaborative filtering network (Liu et al., 2020) has been introduced to aggregate both common and domain-specific user features using Graph Neural Networks (GNNs) that capture high-order user-item connections. Recent studies have further enriched cross-domain recommender systems through the integration of self-supervised learning techniques. For example, C2DSR (Cao et al., 2022) utilizes contrastive learning with both single-domain and cross-domain representations. CCDR (Xie et al., 2022) proposes intra-domain and inter-domain contrastive learning to enhance recommendation performance. Moreover, SITN (Sun et al., 2023) employs self-attention modules as sequence encoders to represent two user-specific sequences from the source and target domains, followed by contrastive learning between them. However, a significant limitation in current cross-domain research is its reliance on correlations (e.g., overlapping users) between source and target data, which constrains its applicability and generalizability. In contrast, text-based zero-shot learning does not has such constraint, allowing for effective transfer even across different datasets. Our proposed EasyRec takes a fundamental step forward in this line of research.
@@
Graph Collaborative Filtering for Recommendation. Graph Neural Networks have recently emerged as a promising approach for recommendation systems, enabling the modeling of complex user-item interactions and capturing high-order dependencies in recommendation (Gao et al., 2023). These models are capable of learning representations of users and items by aggregating their interactions in a graph structure, such as PinSage (Ying et al., 2018), NGCF (Wang et al., 2019), and LightGCN (He et al., 2020). To enhance the representation capacity of GNNs against data sparsity in recommender systems, further studies realize the marriage between self-supervised learning and collaborative filtering with data augmentation. Examples include SGL (Wu et al., 2021), SimGCL (Yu et al., 2022), and HCCF (Xia et al., 2022). These developments showcase the potential of self-supervised graph learning in powering recommenders through graph augmentation based on node self-discrimination.
6. Conclusion
The EasyRec framework effectively integrates LMs to enhance recommendation tasks. Our new paradigm, which is both straightforward and effective, has consistently proven to excel across various scenarios, including text-based zero-shot recommendation and text-enhanced collaborative filtering. At the heart of EasyRec’s success lies an innovative methodology that combines collaborative language model tuning with the transformative capabilities of contrastive learning. This unique approach has empowered EasyRec to capture nuanced semantics and high-order collaborative signals—critical elements that have been instrumental in driving remarkable improvements in recommendation performance. Our extensive experiments, which span a diverse array of datasets, have consistently validated the superiority of EasyRec over existing language models. This robust and generalized performance underscores the framework’s remarkable capacity to adapt to dynamic user preferences, making it well-suited for real-world industry scenarios. Furthermore, the consistent improvements observed across different settings indicate that EasyRec is not only effective but also versatile in its application. Looking ahead, the potential for EasyRec to be seamlessly integrated with multi-modal information presents an enticing frontier for our future investigations.
Appendix A Appendix
@@
Figure 6. An example of large language models-based user profile diversification.
@@
Figure 7. An example of large language models-based item profile diversification.
A.1. Implementation and Training Details
@@
We implemented our EasyRec and conducted all experiments using PyTorch (Paszke et al., 2019). For the transformer-based encoder backbone, we adopted the architecture of RoBERTa (Liu et al., 2019) and utilized its pre-trained parameters as initialization. We trained three versions of EasyRec with varying parameter sizes (small, base, and large), as detailed in Table 2. For the loss function, we set the hyperparameters $\tau$ to 0.05 and $\lambda$ to 0.1. The token masking ratio for masked language modeling is 0.15, and the learning rate is set to $5\times 10^{-5}$. We train the model for 25 epochs. For profile augmentation, we set the diversification time t for LLM-based methods to 3, resulting in 4 user profiles and 5 item profiles per dataset. During training, we evaluate the model every 1000 steps and use the validation interactions from each training dataset to select the optimal model parameters, employing the Recall@20 metric. Detailed implementation of our model is provided in our anonymous released code.
A.2. Datasets and User/Item Profiles

### v4

@@ 4.2 Performance Comparision for Text-based Recommendation (RQ1)
4.2.1 Baseline Methods and Settings
We have included the following state-of-the-art language models as text embedders for comparative evaluation: (i) General Language Models: BERT Devlin et al. (2018), RoBERTa Liu et al. (2019) and BART Lewis et al. (2019); (ii) Language Models for Dense Retrieval: SimCSE Gao et al. (2021), GTR Ni et al. (2021) and BGE Xiao et al. (2023); (iii) Pre-trained Language Models for Recommendation: BLaIR Hou et al. (2024a). We also compare with SoTA text embedding models provided by OpenAI. Details of the baseline models are described in Appendix G.
4.2.2 Result Analysis
The overall comparison of different models is presented in Table 3. This evaluation reveals several observations, which are outlined below:
  • Superiority across Diverse Datasets. Our evaluation consistently shows that the EasyRec outperforms all other models across the datasets spanning different platforms. This provides strong evidence for the effectiveness of the EasyRec. We attribute these improvements to two key factors: i) By injecting collaborative signals into the LMs, we effectively optimized our EasyRec using supervised contrastive learning within the recommendation context. This approach allows the model to inherently encode user and item text embeddings that are well-suited for recommendation tasks. ii) By integrating a diverse array of datasets across multiple categories and utilizing data augmentation techniques to enrich the text descriptions for training, our EasyRec exhibits impressive generalization capabilities, enabling it to effectively handle unseen data.
  • Scaling Law Investigation of EasyRec Model. Our experiments reveal that as the size of the EasyRec increases, its performance consistently improves across all the datasets. This observation reflects a scaling law, where the model’s performance growth is directly correlated with its size. Furthermore, this finding effectively reinforces the validity of text-based recommendation systems. It also validates our approach to training the LMs, which enables it to learn collaborative signals from a new perspective.
4.3 Performance of Text-enhanced CF (RQ2)
Table 4: Recommendation performance in text-enhanced CF. The experiment was conducted on the Steam dataset with 5-runs to obtain the mean results.
@@ 4.3 Performance of Text-enhanced CF (RQ2)
BERT 0.0822 0.1313 0.0663 0.0829
RoBERTa 0.0848 0.1351 0.0684 0.0854
BART 0.0874 0.1383 0.0701 0.0874
SimCSE 0.0877 0.1395 0.0706 0.0881
BLaIR 0.0880 0.1392 0.0708 0.0882
GTR 0.0873 0.1387 0.0706 0.0880
BGE 0.0875 0.1393 0.0705 0.0881
EasyRec 0.0881 0.1402 0.0712 0.0888
@@
BERT 0.0849 0.1347 0.0684 0.0852
RoBERTa 0.0867 0.1374 0.0699 0.0870
BART 0.0887 0.1407 0.0715 0.0891
SimCSE 0.0891 0.1417 0.0719 0.0898
BLaIR 0.0897 0.1418 0.0724 0.0901
GTR 0.0894 0.1417 0.0719 0.0896
BGE 0.0891 0.1407 0.0718 0.0893
EasyRec 0.0908 0.1430 0.0732 0.0908
In addition to our investigation of zero-shot recommendation, we explore the potential of EasyRec as an enhancement when integrated with CF models. To assess the effectiveness of various LMs in CF, we employ two widely used ID-based methods as backbone models: GCCF Chen et al. (2020) and LightGCN He et al. (2020), which were chosen for their proven effectiveness and efficiency. Furthermore, we utilize the advanced model-agnostic text-enhanced framework RLMRec Ren et al. (2024) with contrastive alignment to conduct our investigation. We compare the large versions of both EasyRec and other open-source LMs. The findings from the results in Table 4 are:
  • Compared to backbones, the integration of the text-enhanced framework generally improves the performance for both GCCF and LightGCN. This observation highlights the significance of incorporating text modality (i.e., user/item profiles) into the recommendation paradigm.
  • Among the various LMs, EasyRec consistently achieves the highest performance in the text-enhanced recommenders. This outcome not only illustrates the efficacy of EasyRec for recommendation, but also emphasizes the advantages of incorporating collaborative information into LMs.
4.4 Efficacy of Profile Diversification (RQ3)
@@
Figure 4: Performance w.r.t. data size. "Augmentation Count" indicates the number $t$ of diversified profiles.
@@
  • Effectiveness of Profile Diversification. The increase in the number of diversified profiles enhances model performance, particularly for larger models. This finding underscores the effectiveness of our augmentation approach using LLMs for profile diversification, and emphasizes the significance of increasing training corpus for improved recommendation outcomes.
  • Scaling Relationship: The scaling experiments on both model size and data size reveal a crucial relationship that influences model performance. This demonstrates that our approach of training the LM with collaborative signals follows a scaling law, indicating that model performance benefits from both increased capacity and data volume. Such findings are vital as they provide insights into how model capacity and data availability interact, guiding future research and development.
Moreover, we find that profile diversification can improve recommendation diversity. We conduct experiments and discussions in Appendix F.5.
4.5 Fast Adaptation Case Study (RQ4)
Figure 5: Case study on handling user preference shift.
@@ 4.5 Fast Adaptation Case Study (RQ4)
We visualize all the encoded embeddings using t-SNE Van der Maaten and Hinton (2008), as illustrated in Figure 5, which reveals a significant shift in the user embedding within the feature space. Correspondingly, recommended items transition from basketball-related products to swimming gear, reflecting the user’s changing preferences. Notably, this adjustment is accomplished solely by modifying the user’s profile, without further training of the model. This underscores the efficiency and flexibility in adapting to evolving user preferences.
5 Related Work
LM-Powered Recommender Systems. Recent recommendation methods increasingly incorporate text modalities Yuan et al. (2023); Wu et al. (2024); Zhang et al. (2023), using LM-derived semantic embeddings to boost CTR prediction and sequence recommendation Xi et al. (2023); Geng et al. (2024); Hou et al. (2022); Sheng et al. (2024). Some approaches use text-based agents to boost performance Zhang et al. (2024a, b), with notable examples like ZESRec Ding et al. (2021) and MoRec Yuan et al. (2023) leveraging text embeddings for inductive performance. RLMRec applies information-theoretic text augmentation to ID-based recommenders Ren et al. (2024). However, most systems depend on general-purpose encoders (e.g., BERT) rather than recommendation-tailored LMs. Inspired by BLaIR’s use of feedback and metadata Hou et al. (2024a), we train a recommendation-specific LM on user profiles and collaborative signals to improve zero-shot and text-augmented performance.
@@ 5 Related Work
Cross-Domain Recommendation. Cross-domain recommendation enhances recommendations in one domain by leveraging data from another domain to combat data sparsity and improve personalization Zang et al. (2022). Techniques like graph collaborative filtering Liu et al. (2020) use Graph Neural Networks (GNNs) to aggregate common and domain-specific user features. Recent studies have integrated self-supervised learning, such as C2DSR Cao et al. (2022) with contrastive learning for both single and cross-domain representations, and CCDR Xie et al. (2022) with intra- and inter-domain contrastive learning. SITN Sun et al. (2023) utilizes self-attention to represent user sequences from source and target domains, followed by contrastive learning. However, current cross-domain approaches often rely on correlations between source and target data, limiting their application to zero-shot tasks. In contrast, our proposed EasyRec employs text-based zero-shot learning, removing these constraints and enabling effective transfer across different domains.
@@
Graph Collaborative Filtering for Recommendation. Graph Neural Networks (GNNs) are effective for recommendation by modeling user–item interactions and high-order dependencies Gao et al. (2023). Representative models include PinSage Ying et al. (2018), NGCF Wang et al. (2019) and LightGCN He et al. (2020), which learn node representations via neighborhood aggregation. To address data sparsity, recent studies combine self-supervised learning with collaborative filtering using graph augmentations; notable examples are SGL Wu et al. (2021), SimGCL Yu et al. (2022) and HCCF Xia et al. (2022). These works demonstrate the promise of self-supervised graph learning for recommender systems.
6 Conclusion
This paper presents EasyRec  which integrates LMs to enhance recommendation. Our approach is simple yet effective, excelling in scenarios like text-based zero-shot recommendations and text-enhanced CF. Central to EasyRec’s success is its combination of collaborative LM tuning and contrastive learning, which captures nuanced semantics and high-order collaborative signals, leading to improved recommendations. Extensive experiments demonstrate EasyRec’s superiority over existing models, showcasing its adaptability to changing user preferences and real-world applications.
Acknowledgments
This work is supported by the National Natural Science Foundation of China under Grants 624B2122.
Limitation
While EasyRec shows promising advancements in generalization for zero-shot recommendation, it faces challenges related to data modality diversity. Currently, our approach relies primarily on textual data, whereas items can encompass a richer variety of modalities. The absence of visual inputs, such as images and videos, limits the contextual information we can leverage. These visual elements have the potential to convey aesthetic preferences, cultural trends, and emotional responses, capturing nuances that textual data might overlook. They can also reveal visual patterns linked to user behavior and preferences, including color schemes, styles, and settings, which are essential for effective personalization. Acknowledging these aspects, future work could explore the integration of multimodal data processing techniques, potentially enhancing predictive accuracy and improving the system’s ability to generalize recommendations.
Appendix A Discussion on EasyRec
A.1 Motivation and Technical Positioning
The motivation behind our proposed EasyRec stems from the limitations of existing recommendation algorithms that primarily rely on ID-based representations. While ID-based representations can enhance algorithm performance, they fundamentally restrict the generalization capabilities of recommendation systems. Specifically, models trained in this manner cannot be directly transferred to new datasets in zero-shot recommendation scenarios, where there is no overlap between users and items. Recent studies Ren et al. (2024); Sheng et al. (2024) have shown a growing interest in incorporating language representations (i.e., text embeddings) to replace or complement ID-based representations. This approach aims to improve the generalization and overall performance of these algorithms. However, the predominant reliance on general-purpose language models as text encoders has created a gap in the development of language models specifically tailored for recommendation scenarios, which could yield higher-quality user and item embeddings. To address this gap, we propose utilizing user and item profiles as fundamental textual representations. By training a language model specifically designed for recommendation algorithms, based on collaborative filtering signals and profile diversification, we can generate embeddings that are more suitable for recommendation contexts. This approach promises to deliver improved performance compared to general LMs.
A.2 Relationship between EasyRec and LLM-CF Methods
The proposed EasyRec constitutes a specialized class of language models meticulously trained for recommendation scenarios. Functioning as a textual encoder, its primary objective lies in transforming descriptive textual profiles of users and items into discriminative feature representations compatible with collaborative filtering architectures. The training process employs collaborative filtering signals as self-supervised signals to guide model optimization. The motivation of EasyRec stands distinct from approaches that combine LLMs with CF models as end-to-end frameworks Ren et al. (2024); Yang et al. (2024). By contrast, EasyRec adopts a foundational modeling perspective, aiming to enhance existing text-enhanced algorithms through improved embeddings.
Recently, LLM-CF integration methods predominantly utilize text embeddings as auxiliary features. As empirically validated in our experiments on Text-enhanced Collaborative Filtering (Section 4.3), EasyRec shows superior performance compared to open-source text embedders.
A.3 Can Text-Only Frameworks Outperform ID-Based Recommendation?
ID-based recommendation systems have long served as the predominant paradigm in both industrial deployments and academic research. However, their inherent limitations in cross-domain generalization have recently propelled text-only recommendation approaches – the core focus of EasyRec – to the forefront of research attention. While text-based methods demonstrate enhanced generalization capabilities, the performance hierarchy between text-based and ID-based approaches remains empirically underexplored.
To investigate this research question, we integrate EasyRec with the state-of-the-art text-based framework AlphaRec Sheng et al. (2024), conducting comparative analyses against the ID-based LightGCN backbone. To ensure robustness, we perform multiple runs and report averaged results. The experimental results detailed in Table 5 indicate that the text-only framework provides superior performance compared to ID-based methods.
Table 5: Performance comparison (Steam dataset) between text-based and ID-based recommenders
Model R@10 R@20 N@10 N@20
LightGCN (ID-based) 0.0851 0.1349 0.0686 0.0854
AlphaRec + EasyRec 0.0863 0.1358 0.0689 0.0857
@@ A.3 Can Text-Only Frameworks Outperform ID-Based Recommendation?
Figure 6: An example of large language models-based user profile diversification.
@@
Figure 7: An example of large language models-based item profile diversification.
Appendix B Implementation and Training DetaILD
@@
We implemented our EasyRec and conducted all experiments using PyTorch Paszke et al. (2019). For the transformer-based encoder backbone, we adopted the architecture of RoBERTa Liu et al. (2019) and utilized its pre-trained parameters as initialization. We trained three versions of EasyRec with varying parameter sizes (small, base, and large), as detailed in Table 2. For the loss function, we set the hyperparameters $\tau$ to 0.05 and $\lambda$ to 0.1. The token masking ratio for masked language modeling is 0.15, and the learning rate is set to $5\times 10^{-5}$. We train the model for 25 epochs. For profile augmentation, we set the diversification time $t$ for LLM-based methods to 3. During training, we evaluate the model every 1000 steps and use the validation interactions from each training dataset to select the optimal model parameters, employing the Recall@20 metric. Detailed implementation of our model is provided in the code.
Appendix C Comparison with Zero-shot Frameworks
The exploration of zero-shot recommendation has garnered significant attention, particularly in sequential recommendation scenarios Ding et al. (2021). In our work, we propose a novel zero-shot research line for collaborative filtering (CF) based on semantic profile matching. To demonstrate its effectiveness, we compare it with another fundamental zero-shot framework, LLMRank Hou et al. (2024b), which leverages LLMs for item ranking. Specifically, we implement LLMRank with GPT-4o-mini to rank candidate items based on historical interactions and item titles, while our method employs user/item profile embeddings for feature-based ranking. Following the same experimental protocol, for each user in the test set we construct a candidate set containing 19 negative items and 1 ground-truth item. As shown in Table 6, our approach demonstrates better recall performance while achieving higher efficiency.
Table 6: Performance comparison with other zero-shot frameworks on sports dataset.
Model R@5 R@10 Inference Time
LLMRank 0.6155 0.7822 $\sim$1 hours
EasyRec 0.6332 0.8078 3-4 minutes
Appendix D Profile Generation with Different Large Language Models
To explore whether the trained EasyRec can generalize to different LLMs that under different tokenizers, we regenerate alternative user/item profiles using DeepSeek-V3 Liu et al. (2024) on the Sports dataset and performed a performance validation based on EasyRec-Large. The results on Sports dataset are shown in Table 7, from which we can observed that the profiles generated based on GPT-3.5 and those generated by DeepSeek have a minimal impact on performance. This is because the final profiles are presented in text form. During the training of EasyRec, we exposed it to a large diversity of recommendation-related texts and utilized multiple datasets for joint optimization, effectively avoiding overfitting on specific profile patterns.
Table 7: Performance comparison with various LLM-generated profiles on sports dataset.
Profile Generator R@10 R@20 N@10 N@20
GPT-3.5 0.0423 0.0586 0.0250 0.0294
DeepSeek-V3 0.0417 0.0555 0.0253 0.0290
Appendix E Impact of Training Objectives
To evaluate the impact of different training objectives on the LM’s learning, we also implemented EasyRec-Large training with BPR loss (i.e., one negative item per training sample) for comparison with the contrastive learning results. As shown in Table 8, the model performance generally outperforms that trained with BPR loss, demonstrating the effectiveness of using contrastive learning to incorporate collaborative information into the LMs.
Table 8: Comparison of EasyRec with different learning objectives (where "Contrast" stands for contrastive).
Objective Sports Yelp
R@10 N@10 R@10 N@10
BPR Loss 0.0381 0.0226 0.0028 0.0021
Contrast Loss 0.0396 0.0236 0.0034 0.0026
Appendix F Datasets and User/Item Profiles